OpenStudy (anonymous):

What is the solution to the following equation? 6(x – 3) = 2(13 – x) x= 1 f(x)x^2+5x+6 k(x)4x^2+11x+6 x=11

OpenStudy (anonymous):

I got x=11/2

OpenStudy (abb0t):

It's probably A or D. bcause you can't get \(x^2\)

OpenStudy (anonymous):

6X-18=26-2X 6X+2X=26+18 8X=44 X=11/2

OpenStudy (anonymous):

I got D

OpenStudy (anonymous):

IT is D